# Spectrality of Prime Size Tiles

### v1

@@ 1 Introduction
Fuglede conjectured that being spectral and being tiling are actually equivalent [6], this has been studied extensively and disproved in general when the dimension is at least three [5, 10, 11, 17, 20] but remains open for one dimensional and two dimensional cases. Researches over the Fuglede conjecture in finite Abelian groups as well as in $p$-adic fields have been active in the past years, see e.g., [3, 4, 9, 12, 13, 15, 16, 21] (the full list is of course too long to be enumerated here, this is only an attemp to mention some of the results in one dimensional or two dimensional cases, see perhaps references therein for more on the history and development of this problem).
@@
Here $p$ points being in general linear positions means they are not contained in any hyperplane or affine hyperplane of dimension p2p-2, alternatively this means if we fix any one of these $p$ points, then the p1p-1 vectors formed from the fixed point to the other p1p-1 points are linearly independent in $\mathbb{Z}^{d}$. It is also worth mentioning that for $d\geq p$ the second theorem is acutually already included in [17, Corollary 2.4].
@@ 2 Preliminaries
where $\langle\cdot,\cdot\rangle$ is the Euclidean inner product. It is straigtforward to verify that
@@
Assume the contrary that $\mathbf{1}_{A}$ is not a scalar multiple of $\mathbf{1}_{\mathbb{Z}_{n}^{d}}$, then let $m$ be the smallest multipicity of elements in $A$, i.e.,
@@
The way that Lemma 2 is stated here is slightly different from its original version in [19, Theorem 17], but it is clear that if $\{(t_{1},0,\ldots,0),\ldots,(0,\ldots,0,t_{d})\}$ is a set of periods of $B$, then the least common multiple of $t_{1},\ldots,t_{d}$ can be taken as $n$, and $B^{\prime}$ can be obtained by removing all duplications from $\pi_{n}(B)$.
@@
Let $E$ be an equaivlence class under $\sim$ in $\mathbb{Z}_{n}^{d}$ and $A\subseteq\mathbb{Z}_{n}^{d}$. If $A$ annihilates one element of $E$, then it annihilates entire $E$.
@@
Let $E$ be an equaivlence class under $\sim$ in some finite Abelian group. If $p$ is the smallest prime divisor of $\operatorname{ord}(E)$, then $E$ admits a derived set of size $p$.
@@ 3 Main results
Now as $\pi_{p^{k}}(B^{\prime})$ is a multiset on $\mathbb{Z}_{p^{k}}^{d}$, Lemma 1 indicates that $\mathbf{1}_{\pi_{p^{k}}(B^{\prime})}$ is just a scalar multiple of $\mathbf{1}_{\mathbb{Z}_{p^{k}}^{d}}$, which means $|Z_{p^{k}}^{d}|$ divides $|\mathbf{1}_{\pi_{p^{k}}(B^{\prime})}|$. Since $|\mathbf{1}_{\pi_{p^{k}}(B^{\prime})}|=|B^{\prime}|$, this further indicates that $p^{kd}$ divides $|B^{\prime}|$, consequently we would have $p^{kd}m^{d}=|\mathbb{Z}_{n}^{d}|=|A|\cdot|B|$ is divisible by $p^{kd+1}$, which is a contradiction. ∎
@@
Given any $n$ linearly indpenedent vectors $v_{1},\ldots,v_{n}$ on $\mathbb{Z}^{n}$, there exists a group automorphism of the additive group $\mathbb{Z}^{n}$ that maps these vectors to $(1,0,\ldots,0)$, $(0,1,0,\ldots,0)$, $\ldots$, $(0,\ldots,0,1)$. In particular, if $p$ is a prime number and n=p1n=p-1, then there exists a group automorphism of the additive group $\mathbb{Z}^{p-1}$ that maps these vectors to $(1,0,\ldots,0)$, $(0,2,0,\ldots,0)$, $\ldots$, $(0,\ldots,0,p-1)$.
@@
If n=p1n=p-1, then $\operatorname{diag}(1,2,\ldots,{p-1})$ is also a group autophism on $\mathbb{Z}^{p-1}$, thus the desired automorphism is constructed by a further composition:
$\mathbb{Z}^{p-1}\xrightarrow{\qquad\qquad\qquad V^{\prime}\qquad\qquad\qquad}D\mathbb{Z}^{p-1}\xrightarrow{\qquad\qquad x\mapsto x/D\qquad\qquad}\mathbb{Z}^{p-1}\xrightarrow{\qquad\operatorname{diag}(1,2,\ldots,{p-1})\qquad}\mathbb{Z}^{p-1}.$
@@
Since $p$ is a prime number, the (p1)(p-1)-st roots of unity in the finite field $\mathbb{F}_{p}$ is precisely $1,\ldots,p-1$, hence all entries of the $(p-1)\times(p-1)$ Fourier matrix (i.e., the Vandermonde matrix generated by these roots of unity $1,\ldots,p-1$) are in $\mathbb{Z}_{p}$, and its columns, denoted by $u_{1},\ldots,u_{p-1}$ is a basis (the Fourier basis) on $\mathbb{Z}_{p}^{p-1}$ (understood as a vector space over $\mathbb{F}_{p}$). In particular, since $x$ is a column of it we may without loss of generality take $u_{1}=x$.
@@
The condition that $p$ points have to be in general linear positions is tight in the sense that $3$ colinear points need not be tiling, for example, $\{0,3,4\}$ does not tile $\mathbb{Z}$ (corresponds to $p=3$, and d=1<2=31d=1<2=3-1). The author does not know non-tiling examples for other values of $p$, but would expect that they exist for every pair $p,d$ with d<p1d<p-1.

### v3

@@ 1 Introduction
Fuglede conjectured that being spectral and being tiling are actually equivalent [6], this has been studied extensively and disproved in general when the dimension is at least three [5, 10, 11, 17, 21] but remains open for one dimensional and two dimensional cases. Researches over the Fuglede conjecture in finite Abelian groups as well as in $p$-adic fields have been active in the past years, see e.g., [3, 4, 9, 12, 13, 15, 16, 22] (the full list is of course too long to be enumerated here, this is only an attempt to mention some of the results in one dimensional or two dimensional cases, see perhaps references therein for more on the history and development of this problem).
@@
Here $p$ points being in general linear positions means they are not contained in any hyperplane or affine hyperplane of dimension p2p-2, alternatively this means if we fix any one of these $p$ points, then the p1p-1 vectors formed from the fixed point to the other p1p-1 points are linearly independent in $\mathbb{Z}^{d}$. It is also worth mentioning that for $d\geq p$ the second theorem is actually already included in [17, Corollary 2.4].
@@ 2 Preliminaries
where $\langle\cdot,\cdot\rangle$ is the Euclidean inner product. It is straightforward to verify that
@@
Assume the contrary that $\mathbf{1}_{A}$ is not a scalar multiple of $\mathbf{1}_{\mathbb{Z}_{n}^{d}}$, then let $m$ be the smallest multiplicity of elements in $A$, i.e.,
@@
The way that Lemma 2 is stated here is slightly different from its original version in [20, Theorem 17], but it is clear that if $\{(t_{1},0,\ldots,0),\ldots,(0,\ldots,0,t_{d})\}$ is a set of periods of $B$, then the least common multiple of $t_{1},\ldots,t_{d}$ can be taken as $n$, and $B^{\prime}$ can be obtained by removing all duplications from $\pi_{n}(B)$.
@@
Let $E$ be an equivalence class under $\sim$ in $\mathbb{Z}_{n}^{d}$ and $A\subseteq\mathbb{Z}_{n}^{d}$. If $A$ annihilates one element of $E$, then it annihilates entire $E$.
@@
Let $E$ be an equivalence class under $\sim$ in some finite Abelian group. If $p$ is the smallest prime divisor of $\operatorname{ord}(E)$, then $E$ admits a derived set of size $p$.
@@ 3 Main results
Now as $\pi_{p^{k}}(B^{\prime})$ is a multiset on $\mathbb{Z}_{p^{k}}^{d}$, Lemma 1 indicates that $\mathbf{1}_{\pi_{p^{k}}(B^{\prime})}$ is just a scalar multiple of $\mathbf{1}_{\mathbb{Z}_{p^{k}}^{d}}$, which means $|Z_{p^{k}}^{d}|$ divides $|\pi_{p^{k}}(B^{\prime})|$. Since $|\pi_{p^{k}}(B^{\prime})|=|B^{\prime}|$, this further indicates that $p^{kd}$ divides $|B^{\prime}|$, consequently we would have $p^{kd}m^{d}=|\mathbb{Z}_{n}^{d}|=|A|\cdot|B|$ is divisible by $p^{kd+1}$, which is a contradiction. ∎
@@
Given any $n$ linearly indenpendent vectors $v_{1},\ldots,v_{n}$ on $\mathbb{Z}^{n}$, there exists a group automorphism of the additive group $\mathbb{Z}^{n}$ that maps these vectors to $(1,0,\ldots,0)$, $(0,1,0,\ldots,0)$, $\ldots$, $(0,\ldots,0,1)$. In particular, if $p$ is a prime number and n=p1n=p-1, then there exists a group automorphism of the additive group $\mathbb{Z}^{p-1}$ that maps these vectors to $(1,0,\ldots,0)$, $(0,2,0,\ldots,0)$, $\ldots$, $(0,\ldots,0,p-1)$.
@@
If n=p1n=p-1, then $\operatorname{diag}(1,2,\ldots,{p-1})$ is also a group automorphism on $\mathbb{Z}_{p}^{p-1}$, thus the desired automorphism is constructed by a further composition:
$\mathbb{Z}^{p-1}\xrightarrow{\qquad V^{\prime}\qquad}D\mathbb{Z}^{p-1}\xrightarrow{\;x\mapsto x/D\;}\mathbb{Z}^{p-1}\xrightarrow{\;\operatorname{diag}(1,2,\ldots,{p-1})\otimes I\;}\mathbb{Z}^{p-1}\left(=\mathbb{Z}_{p}^{p-1}\oplus(p\mathbb{Z})^{p-1}\right),$
where $I$ is identity so that $\operatorname{diag}(1,2,\ldots,{p-1})$ only acts on the $\mathbb{Z}_{p}^{p-1}$ part. ∎
@@
Since $p$ is a prime number, the (p1)(p-1)-st roots of unity in the finite field $\mathbb{F}_{p}$ is precisely $1,\ldots,p-1$, hence all entries of the $(p-1)\times(p-1)$ Fourier matrix (i.e., the Vandermonde matrix generated by these roots of unity $1,\ldots,p-1$) are in $\mathbb{Z}_{p}$, and the set of its columns, denoted by $u_{1},\ldots,u_{p-1}$, is a basis (the Fourier basis) on $\mathbb{Z}_{p}^{p-1}$ (understood as a vector space over $\mathbb{F}_{p}$). In particular, since $x$ is a column of it we may without loss of generality take $u_{1}=x$.
@@
The condition that $p$ points have to be in general linear positions is tight in the sense that $3$ colinear points need not be tiling, for example, $\{0,3,4\}$ does not tile $\mathbb{Z}$ (corresponds to $p=3$, and d=1<2=31d=1<2=3-1, see also [18] for a characterization of prime power size tiles in $\mathbb{Z}$). The author does not know non-tiling examples for other values of $p$, but would expect that they exist for every pair $p,d$ with d<p1d<p-1.